# TeleMem: Building Long-Term and Multimodal Memory for Agentic AI

### v2

@@ Abstract
Large language models (LLMs) excel at many NLP tasks but struggle to sustain long-term interactions due to limited attention over extended dialogue histories. Retrieval-augmented generation (RAG) mitigates this issue but lacks reliable mechanisms for updating or refining stored memories, leading to schema-driven hallucinations, inefficient write operations, and minimal support for multimodal reasoning. To address these challenges, we propose TeleMem, a unified long-term and multimodal memory system that maintains coherent user profiles through narrative dynamic extraction, ensuring that only dialogue-grounded information is preserved. TeleMem further introduces a structured writing pipeline that batches, retrieves, clusters, and consolidates memory entries, substantially improving storage efficiency, reducing token usage, and accelerating memory operations. Additionally, a multimodal memory module combined with ReAct-style reasoning equips the system with a closed-loop observe, think, and act process that enables accurate understanding of complex video content in long-term contexts. Experimental results show that TeleMem surpasses the state-of-the-art Mem0 baseline with $19\%$ higher accuracy, $43\%$ fewer tokens, and a $2.1\times$ speedup on the ZH-4O long-term role-play gaming benchmark.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing tasks [5, 37, 29]. However, their effectiveness in long-term interactions remains fundamentally constrained by the context window limitations inherent to Transformer architectures. Even though contemporary long-context models can process hundreds of thousands of tokens [10], simply enlarging the context window does not fully address the issue: as interaction rounds accumulate, models struggle to allocate attention to distant tokens, making it difficult to accurately retrieve user-specific information embedded deep in the dialogue history. As a result, they may overlook key details or even forget important personalized information over time [12], significantly hindering their practical utility in scenarios requiring sustained, long-term relationships with users.
To address the context-window limitations, retrieval-augmented generation (RAG) has become a widely adopted strategy for managing long dialogue histories. RAG systems typically encode past interactions into vector embeddings stored in external datastores and retrieve relevant information via semantic search [21, 44, 8]. Although effective at retrieving information beyond the model’s native context window, traditional RAG pipelines fundamentally lack mechanisms for updating or refining stored memories. Once written into the datastore, memories cannot be reliably modified or deleted, making it difficult to handle evolving user preferences, resolve contradictions, or maintain long-term consistency.
@@
This inherent limitation has motivated the development of more advanced memory architectures capable of structured storage, organization, and dynamic updating of information [40, 36]. For example, MemoryBank [45] introduces a decay mechanism inspired by the Ebbinghaus forgetting curve to selectively preserve salient information. Mem0 [2] proposes a meaning-aware memory framework in which the LLM itself extracts atomic facts, retrieves semantically related memories, and decides whether to add, update, delete, or ignore them in a streaming, turn-by-turn process. Rsum [25] employs hierarchical summarization to distinguish short-term details from long-term abstractions, while Zep [22] represents dialogue history as a temporal knowledge graph to capture causal and temporal dependencies. Collectively, these systems extend RAG by incorporating notions of forgetting, abstraction, and structure, offering more principled approaches for managing long-term conversational memory.
However, despite recent progress, existing memory systems still face three major challenges in real-world deployment. First, maintaining a consistent user profile remains difficult. Current systems rely on dozens of predefined fields, while real conversations typically provide only sparse information. This mismatch forces the model to hallucinate unsupported details and leaves many fields empty, creating unnecessary computational overhead and increasing the likelihood of structural errors. Second, current memory systems exhibit low write efficiency. Each dialogue turn triggers a retrieval step and sends the top-k memories to an LLM for add, delete, update, or no-op decisions, resulting in frequent datastore writes and inefficient API round-trips that degrade throughput and latency. Third, multimodal reasoning capabilities remain limited. Most systems are predominantly text-centric and struggle to integrate or reason over memories derived from images, audio, and other modalities, significantly restricting their use in complex real-world settings.
@@
To tackle these challenges, we propose TeleMem, a unified long-term and multimodal memory system. By extracting only dialogue-supported narrative units, TeleMem avoids schema-driven hallucinations and enables lightweight, precise updates, resulting in a compact and reliable long-term memory. To further address storage and retrieval inefficiencies, TeleMem employs a novel writing pipeline that batches summaries, retrieves related memories, clusters semantically similar entries, and applies an LLM-driven decision step before committing them to persistent storage. This process pre-aggregates and de-duplicates fragmented information, substantially improving throughput while reducing token usage and storage overhead. In addition, TeleMem incorporates a multimodal memory module that transforms raw video streams into event and object memories. Combined with a ReAct-style reasoning framework [38], this design enables a closed-loop observe–think–act process for precise reasoning over complex video content. In summary, our contributions are following:
@@
    We introduce TeleMem, a unified long-term and multimodal memory framework that leverages narrative dynamic extraction to maintain coherent, hallucination-free user profiles.
@@
    We design a structured writing pipeline that batches, retrieves, clusters, and consolidates memory entries, substantially improving storage efficiency, reducing token overhead, and accelerating memory operations.
@@ 2 TeleMem System
TeleMem integrates text-based and multimodal memory into a unified long-term memory system. The text memory module includes profile memory, which captures stable user attributes through lightweight summarization and vector retrieval, and event memory, which models dynamic interaction details using summarization, retrieval, clustering, and consolidation (Sec. 2.1). Complementing these, the multimodal memory module processes video inputs through segmentation, captioning, and cross-modal embedding, and supports video-grounded reasoning via ReAct-style tool use and temporal querying (Sec. 2.2). Together, these components enable TeleMem to maintain consistent, structured, and query-efficient long-term memory across both linguistic and visual modalities.
Figure 1: Overview of the text memory writing pipeline in TeleMem. The system maintains three types of long-term memory: user profile memory $\mathcal{M}_{\mathrm{user}}$, bot profile memory $\mathcal{M}_{\mathrm{bot}}$, and event memory $\mathcal{M}_{e}$. 1). For each dialogue turn $D_{t}$, the model generates role-specific profile summaries $p_{t}^{\mathrm{user}}$ and $p_{t}^{\mathrm{bot}}$, which are written into their respective profile memory stores in parallel. 2). Event memory is constructed through a four-stage batch pipeline. Each turn $D_{t}$ is summarized into one or more textual summaries $s_{i}$, forming the set $\mathcal{S}$. For every summary $s_{i}$, the system retrieves the most relevant event memories from $\mathcal{M}_{e}$, yielding retrieval sets $R_{i}$. All summaries and retrieved items are merged into a unified candidate pool $\mathcal{U}$, which is globally clustered into semantic groups $C_{i}\in\mathcal{C}$. For each cluster $C_{i}$, entries are chronologically ordered and passed to an LLM that assigns an action $A\in\{\textit{add},\textit{delete},\textit{update},\textit{no-op}\}$, producing consolidated event memory updates that are written back into $\mathcal{M}_{e}$.
2.1 Text Memory Module
Overall, the text memory module centers around two complementary operations: memory writing, which transforms incoming dialogue into structured long-term representations through summarization, retrieval, clustering, and refinement (Sec. 2.1.1); and memory reading, which retrieves and integrates relevant profile and event information to support downstream reasoning (Sec. 2.1.2). Together, these operations enable TeleMem to continuously accumulate, organize, and utilize long-term textual knowledge in a coherent and query-efficient manner.
Formally, we denote the dialogue at turn $t$ as $D_{t}$. We maintain three memory stores: a user profile memory $\mathcal{M}_{\mathrm{user}}$, a bot profile memory $\mathcal{M}_{\mathrm{bot}}$ both capturing stable character attributes and relations from their respective perspectives, and an event memory $\mathcal{M}_{e}$, which contains dynamic, temporally grounded interaction summaries.
2.1.1 Memory Writing
The profile and event writing mechanisms enable TeleMem to construct a coherent, dynamically evolving memory space that faithfully reflects both stable role attributes and unfolding narrative developments.
Profile Memory
To capture stable, character-centric information, TeleMem constructs role-specific profile summaries from each dialogue turn $D_{t}$. Because a single turn contains contributions from both interlocutors, the system produces two separate summaries $p_{t}^{\mathrm{user}}$ and $p_{t}^{\mathrm{bot}}$, which are written into the corresponding profile memory stores $\mathcal{M}_{\mathrm{user}}$ and $\mathcal{M}_{\mathrm{bot}}$. Unlike schema-driven methods that rely on predefined fields or incremental slot updates, TeleMem imposes no fixed profile template. Instead, each $p_{t}^{\mathrm{user}}$ and $p_{t}^{\mathrm{bot}}$ is a standalone textual description grounded solely in information explicitly supported by the dialogue. This lightweight formulation avoids schema constraints, prevents unsupported inferences, and keeps profile memory flexible, grounded, and easy to retrieve throughout long-term interactions. Moreover, this role-specific formulation naturally generalizes to settings with additional participants, allowing profile memories to scale seamlessly to multi-party interactions.
Figure 2: Example JSON snippet of a text profile memory entry.
Event Memory
Event memory is responsible for capturing the dynamic, interaction-driven information that emerges throughout a multi-turn dialogue. Different from streaming methods  [2] that update memory turn by turn and often incur redundant retrievals and fragmented writes, our batch formulation enables joint processing of multiple dialogue turns, improving coherence, efficiency, and consolidation quality. Assuming that the system receives a batch of dialogue turns $\{D_{t}\}_{t=1}^{N}$. To transform these into long-term event memory, TeleMem applies a four-stage writing pipeline consisting of summarization, retrieval, clustering, and decision (Fig. 1).
@@ 2.1.1 Memory Writing
  • Summarization (parallel across turns): Each dialogue turn $D_{t}$ is independently summarized into a set of one or more textual summaries, denoted as $S_{t}$. We denote the collection of all summaries produced from a batch of dialogue turns as
    $\mathcal{S}=\bigcup_{t=1}^{N}S_{t},$
@@
  • Retrieval (parallel across summaries): For each summary $s_{i}\in\mathcal{S}$, the system performs vector retrieval from the memory store $\mathcal{M}_{e}$, obtaining the top-$k$ related memories:
    ${R}_{i}=\text{Top-}k\!\left(\text{sim}(s_{i},\mathcal{M}_{e})\right).$
  • Clustering (global, non-parallel): A unified candidate set is constructed as
    $\mathcal{U}=\left(\bigcup_{i=1}^{N}{R}_{i}\right)\cup\mathcal{S},$
    upon which global semantic clustering is applied to obtain clusters
    $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{m}\},\quad C_{i}\subseteq\mathcal{U},$
    where each cluster groups entries describing semantically related events or topics.
  • Decision (parallel across clusters): For each cluster $C_{i}$, the system orders all entries temporally and feeds them to an LLM, which determines for each item whether to perform an action $\mathcal{A}\in\{\emph{add},\emph{delete},\emph{update},\emph{no-op}\}$. This process resolves redundancies and inconsistencies and produces refined memory entries that are subsequently re-embedded and written back into $\mathcal{M}_{e}$.
Figure 3: Example JSON snippet of a text event memory entry.
2.1.2 Memory Reading
Given a query $q$, TeleMem retrieves relevant information from three memory stores: the user profile memory $\mathcal{M}_{\text{user}}$, the bot profile memory $\mathcal{M}_{\text{bot}}$, and the event memory $\mathcal{M}_{e}$. For each store, the embedding of $q$ is used to retrieve the top $5$ most similar entries. These retrieved items provide stable user attributes, stable bot attributes, and dynamic interaction context. They are then inserted into the LLM prompt as external memory, enabling retrieval-augmented reasoning that remains consistent with long-term roles and past events.
2.2 Multimodal Memory Module
Beyond long-term textual inputs, we extend the memory architecture to support multimodal video memory. Our design targets two key requirements of long-horizon video understanding: (1) retaining fine-grained visual grounding while (2) enabling efficient retrieval and reasoning over long sequences. To this end, we build a multimodal memory module that captures both event-level semantics and entity-level information (Sec. 2.2.1), and supports flexible, tool-driven access through an agentic reading mechanism (Sec. 2.2.2).
2.2.1 Memory Writing
We first divide the video stream into 10-second clips and uniformly sample two frames from each clip as the basis for memory extraction. The writing pipeline produces two complementary memory types: event memory and key–value object memory.
Event Memory
For each clip, a VLM generates a concise caption using the sampled frames. These clip-level event memories capture localized semantic events and remain fixed once written, forming a chronological record of the video $\mathcal{M}_{e}$.  [^1] We slightly overload the notation here for simplicity of presentation.
Figure 4: Example JSON snippet of a video event memory entry.
Key–Value Object Memory
For each clip, a VLM generates person and object entities and store them in a key–value structure. Each key specifies stable identity attributes including name, appearance, identity cues, and first occurrence, while each value captures the entity’s behavior or state in the clip. To respect the LLM’s context length constraints, we avoid consolidating all object entries at once. Instead, raw key–value records are accumulated and updated in batches by a text LLM, which incrementally merges redundant information and produces temporally coherent entity summaries. After all batches are processed, we obtain the final consolidated object memory $\mathcal{M}_{\mathrm{obj}}$.
Figure 5: Example JSON snippet of an object memory entry.
2.2.2 Memory Reading
We adopt a ReAct-style agent to read from the multimodal memory, where the agent iteratively selects among three tools to retrieve, refine, or verify information. These tools operate over the text-based memory stores ($\mathcal{M}_{e}$ and $\mathcal{M}_{\mathrm{obj}}$) or the original video clips (Alg. 1).
video.retrieval
This tool performs vector-based retrieval between a textual query and the text-based memory store ($\mathcal{M}_{e}$ and $\mathcal{M}_{\mathrm{obj}}$). It returns the clip indices, specifically the start and end timestamps, that correspond to the most relevant memory entries. This allows the agent to efficiently localize the video segments associated with the query.
video.rag
This tool also conducts vector retrieval against the text-based memory store, but instead of returning clip boundaries, it selects the top-$k$ most relevant memory entries and feeds them into the LLM prompt. The LLM then synthesizes an aggregated, query-grounded response using these retrieved memory elements.
video.qa
This tool performs vision-language question answering over a specified video clip. After the agent identifies a relevant clip, VQA is used to extract fine-grained visual details such as appearance attributes, spatial relations, or object states that may not be available in textual memory.

### v4

@@ Abstract
Large language models (LLMs) achieve strong performance on many NLP tasks but remain limited in long-term interactive settings due to finite context windows and degraded recall over extended histories. Retrieval-augmented generation (RAG) alleviates this bottleneck, yet conventional pipelines treat memories as independent fragments and lack principled mechanisms for consolidation, update, and causal organization, leading to fragmented context and unstable long-horizon reasoning. We propose TeleMem, a unified long-term and multimodal memory system that organizes memory as structured and evolvable semantic trajectories. TeleMem maintains coherent and hallucination-resistant user profiles by extracting only dialogue-grounded narrative units, and amortizes write costs through a structured writing pipeline that batches summaries, retrieves related memories, clusters semantically aligned entries, and performs LLM-based consolidation before persistent storage. At the retrieval level, memories are organized into a threaded directed acyclic graph (DAG), enabling dependency-aware closure-based retrieval that reconstructs coherent causal context for long-horizon reasoning. In addition, TeleMem integrates a multimodal memory module with a ReAct-style reasoning framework to support closed-loop observe–think–act reasoning over complex video content. Extensive experiments demonstrate that TeleMem outperforms the state-of-the-art Mem0 baseline by 19% in accuracy on the ZH-4O benchmark, while reducing token usage by 43% and achieving a $2.1\times$ speedup.
@@ 2 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing tasks [5, 37, 29]. However, their effectiveness in long-term interactive settings remains fundamentally constrained by the finite context window of Transformer architectures. Even though recent long-context models can process hundreds of thousands of tokens [10], simply enlarging the context window does not resolve the core challenge. As interaction histories grow, models struggle to allocate attention to distant yet critical information, leading to degraded recall of user-specific facts and unstable long-horizon reasoning [12]. These limitations hinder the deployment of LLM-based agents in scenarios that require persistent personalization, continuous learning, and traceable decision making.
Beyond capacity, a more fundamental issue lies in how long-term information is represented and organized. Effective memory must preserve not only semantic similarity, but also temporal order, causal dependency, and state evolution across interactions. Without such structure, retrieved contexts easily become fragmented, omit critical prerequisites, and induce inconsistent reasoning.
Retrieval-augmented generation (RAG) has emerged as a practical solution to extend effective memory beyond the native context window by encoding past interactions into vector embeddings and retrieving relevant entries via semantic search [21, 44, 8]. While effective at scaling storage capacity, conventional RAG systems treat memories as independent and unordered fragments. Such designs lack principled mechanisms for updating, consolidating, or deleting previously written memories, making it difficult to accommodate evolving user states, resolve contradictions, or preserve causal consistency over long horizons.
@@
Recent work has therefore explored more structured and adaptive memory architectures that extend RAG with mechanisms for abstraction, forgetting, and dynamic updates [40, 36]. MemoryBank [45] introduces a decay mechanism inspired by the Ebbinghaus forgetting curve to prioritize salient information. Mem0 [2] enables meaning-aware memory operations by allowing an LLM to extract atomic facts and decide whether to add, update, delete, or ignore them in a streaming manner. Rsum [25] applies hierarchical summarization to separate short-term details from long-term abstractions, while Zep [22] organizes dialogue history as a temporal knowledge graph to capture causal dependencies. Although these systems significantly advance long-term memory management, several practical challenges remain.
First, maintaining a consistent and reliable user profile remains difficult. Many systems rely on large predefined schemas, while real conversations provide sparse and noisy signals. This mismatch often induces hallucinated attributes, incomplete fields, and unnecessary structural complexity. Second, write efficiency remains low. Most pipelines invoke retrieval and LLM-based decision making at every dialogue turn, resulting in frequent datastore writes, excessive API calls, and degraded throughput and latency. Third, multimodal reasoning capabilities are still limited. Existing memory systems are predominantly text-centric and struggle to integrate or reason over visual, auditory, and temporal information, restricting their applicability in realistic embodied or multimedia environments.
@@
To address these challenges, we propose TeleMem, a unified long-term and multimodal memory system. TeleMem organizes memory as structured and evolvable semantic trajectories rather than isolated fragments, enabling dependency-aware reasoning and consistent context restoration. It extracts dialogue-supported narrative units to avoid schema-driven hallucinations and maintain compact, reliable user profiles. At the storage level, TeleMem employs a structured writing pipeline that batches summaries, retrieves related memories, clusters semantically aligned entries, and performs LLM-based consolidation before committing them to persistent storage. This design amortizes write costs by reducing per-turn LLM invocations, de-duplicates fragmented information, and substantially improves throughput and token efficiency. At the retrieval level, memories are organized into a threaded directed acyclic graph (DAG), enabling closure-based retrieval that reconstructs coherent causal context for long-horizon reasoning. In addition, TeleMem incorporates a multimodal memory module that converts raw video streams into event and object memories, and integrates a ReAct-style reasoning framework [38] to support closed-loop observe–think–act reasoning over complex visual content.
In summary, our main contributions are:
@@
    We introduce TeleMem, a unified long-term and multimodal memory framework that maintains coherent and hallucination-resistant user profiles through narrative-driven memory extraction and structured organization.
@@
    We propose a structured memory graph with threaded dependencies and closure-based retrieval, enabling dependency-aware context reconstruction and stable long-horizon reasoning.
@@
    We design an efficient memory writing pipeline that performs batched retrieval, semantic clustering, and LLM-based consolidation, substantially reducing storage redundancy, token overhead, and end-to-end latency.
  4. 4.
@@ 3 TeleMem System
Overview.
Long-horizon agents require memory that preserves not only semantic similarity but also the continuity of experience, including temporal order, causal dependency, and state evolution. Conventional retrieval-augmented generation (RAG) systems store memories as independent vector chunks, which often leads to fragmented context, missing prerequisites, and unstable reasoning when interactions span long time horizons. To support stable accumulation and continuous learning, memory must be threaded into coherent trajectories rather than retrieved as isolated fragments.
We address this requirement by organizing long-term memory as a directed acyclic graph (DAG), where memory states are connected through explicit structural relations to form traceable memory threads (Sec. 3.1), upgrading memory from an unordered collection of embeddings to a structured and evolvable substrate that supports causal reasoning and context restoration. The overall architecture tightly couples three functional components: (1) a representation layer that converts raw textual and multimodal interactions into semantic memory states and performs consolidation and updates (Sec. 3.2); (2) a graph layer that incrementally organizes these states into a persistent memory graph; and (3) a memory reading mechanism that performs dependency-aware closure-based retrieval to reconstruct coherent causal context for downstream reasoning (Sec. 3.3). This unified design supports scalable memory growth while preserving structural consistency and low-latency access in practice (Fig. 1).
Figure 1: Comparison between conventional linear memory and the proposed threaded memory with closure-based retrieval. Left: Linear memory retrieves isolated entries based on similarity, resulting in fragmented memories with a lack of causal structure. Right: Threaded memory organizes memory states as a directed acyclic graph (DAG). Given a query seed (gray nodes), closure-based retrieval expands along dependency edges to recover all prerequisite nodes, which are then serialized in chronological order to reconstruct coherent causal context.
3.1 Memory Graph Formulation
We formalize long-term memory as a directed acyclic graph (DAG)
$G=(V,E),$ (1)
where $V$ denotes the set of memory nodes and $E$ denotes directed edges. Each node $v_{i}\in V$ is associated with an effective timestamp $\tau(v_{i})$, which reflects the semantic validity order rather than wall-clock time. All edges satisfy the strict temporal constraint
$\tau(p)<\tau(v_{i}),$ (2)
which guarantees acyclic evolution. To avoid isolated memory fragments and ensure global connectivity, we introduce a virtual root node $v_{0}$ that anchors nodes without explicit predecessors.
3.1.1 The Node: A Unified Semantic Container
Each node represents a stabilized semantic memory state produced by the representation layer and serves as a unified container for long-term knowledge.
Semantic categories.
Nodes may encode heterogeneous semantic states, including: (1) profile states, summarizing stable user and bot attributes (Fig. 3); (2) event states, capturing dynamic interaction-level semantics from textual dialogue or video clips (Fig. 4, Fig. 5); and (3) entity and object states, representing persistent multimodal entities and their temporal evolution (Fig. 6).
Node content.
Each node stores three components: (i) a consolidated semantic content representation $\mathrm{Content}(v_{i})$, (ii) an embedding vector $e_{i}$ for similarity-based retrieval, and (iii) an effective timestamp $\tau(v_{i})$ indicating its semantic validity. This unified abstraction decouples semantic representation from storage topology, enabling consistent indexing, consolidation, and retrieval across modalities.
3.1.2 The Edge: Threaded Causal Skeleton
Edges define how semantic states are connected and evolve over time. A directed edge $p\rightarrow v_{i}$ indicates that node $v_{i}$ is semantically conditioned on prior node $p$. Together with the temporal constraint $\tau(p)<\tau(v_{i})$, the graph enforces monotonic evolution and prevents circular dependencies.
Minimal causal skeleton.
To preserve the threaded semantics and avoid redundant dependencies, we maintain the graph as a minimal causal skeleton, where each edge represents an irreducible dependency rather than a shortcut implied by transitive paths.
Pruning criterion.
An edge $(p\rightarrow v_{i})$ is retained if and only if there exists no alternative directed path from $p$ to $v_{i}$ in the remaining graph:
$(p\rightarrow v_{i})\in E\;\Longleftrightarrow\;p\not\leadsto v_{i}\text{ in }G\setminus\{(p\rightarrow v_{i})\},$
where $a\leadsto b$ denotes reachability via a directed path. Edges violating this criterion are pruned as transitive or redundant dependencies. This effectively maintains a transitive reduction of the dependency DAG. For example, if $pp\rightarrow p$ and $p\rightarrow v_{i}$ exist, then $pp$ can reach $v_{i}$ through $p$, and the shortcut edge $pp\rightarrow v_{i}$ is removed.
Memory threads.
We define a memory thread as any directed path $v_{0}\rightsquigarrow v$ in the DAG, where nodes are strictly ordered by increasing effective timestamps. Such root-to-node paths provide a traceable evolution chain, capturing how topics, states, or beliefs accumulate and change over time as the path progresses chronologically.
By preserving only minimal dependencies under the pruning criterion, the graph avoids transitive shortcuts and keeps these root-to-node threads compact and interpretable. As a result, threads serve as an explicit causal backbone that supports traceable reasoning, dependency-aware retrieval (e.g., closure construction), and stable long-term knowledge accumulation.
3.2 Memory Graph Updating
Long-term memory continuously evolves as new interactions arrive and existing memories are refined. Our system explicitly separates representation updates (how semantic node contents are consolidated) from index updates (how nodes are organized and connected in the graph), and supports both offline batch processing and online incremental maintenance.
Figure 2: Overview of the offline node construction pipeline. A batch of dialogue turns $\{D_{t}\}_{t=1}^{N}$ is first summarized in parallel into textual summaries $s_{i}$. Each summary retrieves related historical memories to form retrieval sets $R_{i}$. All summaries and retrieved candidates are merged into a unified pool $\mathcal{U}$ and globally clustered into semantic groups $C_{i}$. For each cluster, an LLM performs consolidation and assigns an action $\mathcal{A}\in\{\emph{add},\emph{delete},\emph{update},\emph{no-op}\}$ to produce stabilized memory updates. Parallelizable stages across turns, summaries, and clusters are indicated in the diagram.
3.2.1 Graph Update Operators
We define two basic operators for maintaining the graph index: Insert and ReInsert. Both operators enforce the temporal constraint and the pruning criterion (Sec. 3.1.2) and share the same edge construction logic: candidate retrieval, pruning, and edge materialization.
Insert
[…]
  • Candidate retrieval. We restrict the parent search space of $v_{i}$ to historical nodes that satisfy the temporal constraint. The candidate parent set is defined as the Top-$K$ most similar historical nodes:
    $\mathcal{H}(v_{i})=\text{Top-}K\!\Big(\text{sim}\big(v_{i},\;\{v_{j}\in V\mid\tau(v_{j})<\tau(v_{i})\}\big)\Big).$
    This step bounds parent selection to a local semantic neighborhood and avoids quadratic all-pairs comparison.
  • Redundancy pruning. To preserve the threaded causal semantics and avoid transitive or dominated dependencies, we retain only irreducible parents. A candidate parent $p\in\mathcal{H}(v_{i})$ is kept if and only if there exists no alternative path from $p$ to $v_{i}$ through another candidate:
    $\mathcal{P}(v_{i})=\Big\{p\in\mathcal{H}(v_{i})\;\Big|\;\nexists q\in\mathcal{H}(v_{i})\setminus\{p\}\text{ such that }p\leadsto q\Big\},$
    where $p\leadsto q$ denotes reachability in the subgraph induced by historical nodes $\{v_{j}\in V\mid\tau(v_{j})<\tau(v_{i})\}$.
  • Edge materialization. The remaining parents form the incoming edges of $v_{i}$:
    $E_{\mathrm{in}}(v_{i})=\{(p\rightarrow v_{i})\mid p\in\mathcal{P}(v_{i})\}.$
    These edges are added to the graph index.
The Insert operator does not modify any existing node or edge except adding $E_{\mathrm{in}}(v_{i})$, and therefore preserves the global DAG property.
ReInsert.
Given an existing node $v$, ReInsert(v,G)(v,G) re-attaches $v$ to the graph by recomputing its incoming edges using the same procedure as Insert. The operator removes all previous incoming edges of $v$ and materializes a new parent set based on its current representation and the pruning criterion. This local repair is triggered when the semantic content or embedding of $v$ changes, or when its previous dependencies become stale due to upstream updates or deletions. ReInsert updates only the neighborhood of $v$ and preserves the global DAG property and threaded semantics.
3.2.2 Offline Batch Updates
Offline updates are triggered during cold start, large-scale cleanup, or periodic maintenance. They jointly optimize node representations and graph structure at scale.
Node construction.
Historical interactions are processed in batches to construct stabilized and compact semantic node representations. Rather than writing fragmented memories incrementally, the system jointly consolidates multiple dialogue and multimodal segments to reduce redundancy, resolve inconsistencies, and produce high-quality long-term representations.
Let $\mathcal{M}$ denote the current memory store, i.e., the set of existing graph nodes equipped with their embedding index. Assuming that the system receives a batch of dialogue turns $\{D_{t}\}_{t=1}^{N}$, the representation layer applies a four-stage writing pipeline consisting of summarization, retrieval alignment, global clustering, and LLM-based consolidation (Fig. 2):
@@ 3.2.2 Offline Batch Updates
  • Summarization (parallel across turns). Each dialogue turn $D_{t}$ is independently summarized into one or more textual summaries, denoted as $S_{t}$. The collection of all summaries from the batch is
    $\mathcal{S}=\bigcup_{t=1}^{N}S_{t}.$
@@
  • Retrieval alignment (parallel across summaries). For each summary $s_{i}\in\mathcal{S}$, the system performs vector retrieval against the memory store $\mathcal{M}$ to identify the top-$k$ most related existing nodes:
    $R_{i}=\text{Top-}k\!\left(\text{sim}(s_{i},\mathcal{M})\right).$
    This step aligns new content with historical memory to expose semantic redundancy and potential updates.
  • Global semantic clustering (non-parallel). All new summaries and retrieved candidates are merged into a unified candidate pool
    $\mathcal{U}=\left(\bigcup_{i}R_{i}\right)\cup\mathcal{S},$
    which is globally clustered into semantic groups
    $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{m}\},\quad C_{j}\subseteq\mathcal{U}.$
    Each cluster aggregates semantically related content across both new and historical memory.
  • LLM-based consolidation (parallel across clusters). For each cluster $C_{j}$, entries are temporally ordered and passed to an LLM, which determines an action
    $\mathcal{A}\in\{\emph{add},\emph{delete},\emph{update},\emph{no-op}\}$
    for each item. This step resolves redundancy and inconsistency and produces refined semantic representations that are re-embedded and emitted as stabilized graph nodes.
Edge construction.
[…]
All locally constructed edges are aggregated to form the final graph $G=(V,E)$. This parallel construction avoids quadratic all-pairs comparison while preserving temporal validity and the pruning criterion, enabling scalable offline rebuilding.
3.2.3 Online Incremental Updates
Online updates handle real-time interaction with low latency by incrementally updating both (i) the node representation (content/embedding) and (ii) the graph index (incoming edges).
Node incremental update.
In the online setting, node construction follows the same abstraction pipeline as offline consolidation, but operates on a single interaction instance (i.e., batch size =1=1) and omits global clustering for low latency. Given a new dialogue turn or multimodal observation $D_{t}$, the representation layer performs:
  • Summarization. $D_{t}$ is summarized into one or more candidate semantic nodes, capturing distinct facts, events, or state changes expressed in the turn.
  • Retrieval alignment. Each candidate node is matched against the current memory store to identify highly related existing nodes, exposing potential redundancy or update targets.
  • LLM-based decision. Based on the candidate content and retrieved context, an LLM decides whether to add a new node, update an existing node, or perform no-op.
This lightweight pipeline enables real-time semantic state extraction while maintaining consistency with the offline consolidation logic.
Edge incremental update.
Online edge maintenance updates the graph index by invoking the Insert and ReInsert operators (Sec. 3.2.1), while preserving the temporal constraint and the pruning criterion (Sec. 3.1.2).
  • Add. A newly created node $v$ is attached to the graph by invoking Insert(v,G)(v,G). By the temporal constraint, $v$ can only depend on historical nodes; therefore, no existing node needs to be modified.
  • Update. When an existing node $v$ is updated, its semantic representation or embedding may change. We invoke ReInsert(v,G)(v,G) to recompute and replace its incoming edges. To preserve threaded consistency in the downstream structure, all direct children $c\in\mathrm{Ch}(v)$ are additionally refreshed by calling ReInsert(c,G)(c,G). Although these children remain temporally valid, their optimal attachment points under the pruning criterion may change after $v$ is revised.
  • Delete. When a node $v$ is deleted, it is marked as a tombstone and its children $\mathrm{Ch}(v)$ are collected as orphans. Each orphaned child $c$ is repaired by invoking ReInsert(c,G)(c,G), which re-attaches $c$ to alternative valid historical parents. This prevents memory threads from breaking due to the removal of an intermediate dependency.
Remark.
All online updates perform bounded local modifications suitable for low-latency interaction, while periodic offline rebuilding re-optimizes the topology. This process amortizes local approximation errors and preserves long-term structural consistency of the memory graph.
3.3 Memory Reading
Storing memory well is not enough; it must also be retrieved in a way that restores causal context. We therefore replace “Top-$K$ fragment stitching” with closure-based retrieval, which constructs a minimal closed subgraph that contains both the most relevant evidence and its necessary prerequisites.
3.3.1 Retrieval Paradigm: Minimal Closed Subgraph
Given a query $q$, the system retrieves a closed subgraph that contains the most relevant memory threads together with all their required historical dependencies, while avoiding irrelevant sibling branches whenever possible.
Seed identification.
We first retrieve a small set of seed nodes by Top-$K$ embedding similarity between $q$ and graph nodes. If no confident match exists, the virtual root node is used as a fallback seed to preserve global connectivity.
Closure expansion.
Starting from seed nodes, the system recursively traverses dependency edges backward to collect all reachable ancestor nodes until the virtual root node is reached. This produces a closed subgraph in which every node’s prerequisites are included. In practice, if the closure grows excessively large, the traversal can be bounded by a maximum depth or guided by lightweight heuristics (e.g., relevance filtering or budgeted expansion) to control latency and context size.
Context linearization.
All nodes in the closure are sorted by timestamp and serialized into a linear context sequence, optionally augmented with lightweight structural markers, and injected into the LLM prompt for downstream reasoning.
By reconstructing complete causal threads rather than sampling isolated memory fragments, closure-based retrieval preserves prerequisite consistency, reduces thread confusion, and improves long-horizon reasoning stability.
3.3.2 Reasoning Paradigm: ReAct-Style Multimodal Agent
For complex multimodal queries, memory reading is performed by a ReAct-style agent that iteratively executes think–act–observe cycles over textual memory and raw video content (Alg. 1).
Reasoning.
Based on the query and the accumulated interaction history, the agent reasons about what information is missing, which hypotheses require verification, and whether additional evidence is needed from memory or video.
Action.
The agent selects among three tools to refine its knowledge:
  • video.retrieval: retrieve relevant clip timestamps by vector similarity over textual memory;
  • video.rag: aggregate top-$k$ textual memory entries into the LLM prompt for grounded summarization;
  • video.qa: perform vision-language question answering over a specified video clip to extract fine-grained visual details.
Observation.
The outputs returned by the selected tools (retrieved memory entries, localized clips, or visual answers) are appended to the agent history and become the observations for the next reasoning step.
Through iterative think–act–observe cycles, the agent progressively refines its understanding of the multimodal context and produces a grounded response.